For the 3 year ending December 31,2014 through December 31, 2016. a corporation combines a federal and provincial tax payable is as follows:in 2014 the net income tax was $153640, in 2015 it was $186540 and in 2016 its expected to be $186540.Based on this information, calculate the amount of instalment if any is required for the following cases:

1


(a) small corporation(small CPCC)

(b) large corporation(large CPCC)

(c) small corporation assume the combined federal and provincial taxes payable for the year ending December 31 2015 were $163420 instead of $186540. (18 ma




1
Expert's answer
2020-10-05T14:19:50-0400

A) the net income tax was $153640,


it was $186540 

186540- 153640= 32900


B) the net income tax was $ 153640


186540-153640 =32900


C) small corporation assume the combined federal and provincial taxes payable for the year ending December 31 2015 were $163420 instead of $186540. (18 ma

163420-153640 =7220